OpenStudy (anonymous):

when trying to Find the sum of this infinite geometric series. what is r? 12+4+4/3+...

OpenStudy (anonymous):

1/3=r

OpenStudy (anonymous):

\[r=\frac{ t2 }{ t1 } = \frac{ 4 }{ 12 } = \frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

\[Sum=\frac{ a }{ 1-r }\] \[r<1\]

OpenStudy (anonymous):

lol \[\frac{4}{12}=\frac{4/3}{4}=...=r\]

OpenStudy (anonymous):

009infi is rt